# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ IV.2 Triangular lattice (spin-${1\over 2}$): $G_{b}=p3\times{Z}_{2}^{\mathcal{T}}$
one obtains the full symmetry group $F^{c}_{f}$. From the band representation, $\mathcal{C}$-minimal zero modes are found at the $\Gamma$ point and the three M points.
@@
$D(C_{3}^{z})=-I_{4},\ \ \ D(\mathcal{T})K=-\mathrm{i}I_{2}\otimes\sigma_{y}K,$
in the bases $\Psi^{\dagger}_{k}$. Without $\mathcal{C}$, the little co-group becomes $F_{f}(k)/Z_{2}^{f}=\mathscr{C}_{3}\times Z_{2}^{\mathcal{T}}\times Z_{2}^{\mathcal{P}}$. Defining $\tilde{\mathcal{P}}=\mathcal{T}\mathcal{P}$, then the above rep correspond to a set of minimal nonzero modes of the case (6) in the Tab.1. When $\mathcal{C}$ is turned on, then the rep reduces to two sets of irreducible zero modes which cannot couple to each other. Furthermore, the $k\cdot p$ theory indicates that the lowest order dispersion is cubic (see Sec.III.1), and around the $\Gamma$ point, the third-order $k\cdot p$ Hamiltonian can be written as
@@
For the same reason as the $\Gamma$ point, the $\mathcal{C}$ symmetry turns the minimal nonzero modes into two sets of uncoupled zero modes. The $k\cdot p$ theory indicates that the energy dispersion is linear at the $M$ points (see Sec.III.1), for instance, the $k\cdot p$ Hamiltonian for the $(0,{1\over 2})$ point is given by
@@ IV.3 Triangular lattice (spinless): $G_{b}=p6m\times Z_{2}^{\mathcal{T}}$
The band representation of the above group $F^{c}_{f}$ indicates that the high symmetry lines on the mirror planes $\mathcal{M}_{1,2,3}$ (i.e. the lines linking $\Gamma$ and $K,K^{\prime}$ points) are nodal lines with 2-fold zero modes.
@@
(a) band structure with $\mathcal{M}^{\prime}_{1,2,3}$
(b) zero modes inf (a)
(c) band structure without $\mathcal{M}^{\prime}_{1,2,3}$
(d) zero modes in (c)
@@ IV.4 Multi-node $Z_{2}$ QSL on Honeycomb lattice: $G_{b}=p\bar{3}1m\times Z_{2}^{\mathcal{T}}$
$\mathcal{T}^{\prime}=(P_{f}^{B}e^{-\mathrm{i}{\tau_{y}\over 2}{\pi}}e^{-\mathrm{i}{\sigma_{y}\over 2}{\pi}}K||\mathcal{T}),\ \mathcal{I}^{\prime}=(P_{f}^{B}||\mathcal{I}),$
@@
The spectrum of the Hamiltonian (31) can support a series of zero modes in the BZ and the corresponding state is a multi-node QSL. For instance, Fig.11 shows the band structure with parameters $\eta_{0}=2,\eta_{1}=1.2,\eta_{2}=0.8,\eta_{3}=0.5,\eta_{4}=0.2,\eta_{5}=1.2$ and $\eta^{\prime}_{0}=1.6,\eta^{\prime}_{1}=0.2,\eta^{\prime}_{2}=0.8,\eta^{\prime}_{3}=2.5,\eta^{\prime}_{4}=0,\eta^{\prime}_{5}=0$. There are totally $20$ cones, where the $2$ cones in green color are irreducible zero modes protected by the little co-group $F_{f}(K)=\mathscr{C}^{\prime}_{3v}\times Z_{2}^{\mathcal{I}^{\prime}\mathcal{T}^{\prime}}$; the $6$ cones in blue color are protected by crossing of energy bands carrying different quantum numbers of the mirror symmetries $\mathcal{M}^{\prime}_{1,2,3}$; the $12$ cones in red color are protected by $\pi$-quantized Berry phase due to the $\mathcal{I^{\prime}T^{\prime}}$ symmetry.
@@
Figure 11: (a)The Band structure of the Hamiltonian (31) with parameters given in the maintext. The number of the cones is ‘$12+6+2$’ where the ‘$12$’ is protected by the effective $\mathcal{I^{\prime}T^{\prime}}$ symmetry, the ‘$6$’ is protected by the effective mirror reflection symmetries $\mathcal{M}^{\prime}_{1,2,3}$, and the ‘$2$’ form a set of irreducible zero modes; (b) the dots illustrate the positions of the zero modes in (a).
The two high symmetry points $K=({1\over 3},{2\over 3})$ and $K^{\prime}=({2\over 3},{1\over 3})$ have the little co-group $F_{f}(k)/Z_{2}^{f}=\mathscr{C}^{\prime}_{3v}\times Z_{2}^{\mathcal{I^{\prime}T^{\prime}}}\times Z_{2}^{\mathcal{PT^{\prime}}}$. All of the $C^{\prime}_{3c}$, $\mathcal{M^{\prime}}_{1,2,3}$, $\mathcal{I}^{\prime}\mathcal{T}^{\prime}$ symmetries can be treated as effective charge conjugation, and the two fold irreducible zero modes are represented as
@@
According to Eq.(III.1) one has $a_{({\rm v\times ph})}=1$. So the dispersion around the $K$ and $K^{\prime}$ points is linear and the $k\cdot p$ Hamiltonian has one free parameter $v$, yielding
$H_{\rm eff1}=v(\delta k_{x}\mu_{x}+\delta k_{y}\mu_{y}).$
Similarly, the $k\cdot p$ Hamiltonian for the cones at the high symmetry lines with mirror symmetries $\mathcal{M}^{\prime}_{1,2,3}$ have two free parameters $u,v$. For example, around the left blue cone, the effective Hamiltonian reads
$H_{\rm eff2}=$ $u\delta k_{x}{\sqrt{2}\over 2}(\cos\theta_{2}\mu_{x}-\sin\theta_{2}\mu_{y})$
$+v\delta k_{y}{\sqrt{2}\over 2}(\sin\theta_{2}\mu_{x}+\cos\theta_{2}\mu_{y})$
@@
$H_{\rm eff}=(e\mu_{x}+f\mu_{y})\delta k_{x}+(h\mu_{x}+m\mu_{y})\delta k_{y},$ (32)
which has four parameters $e,f,h,m$.
@@ V Discussion and Conclusion
To be Majorana or not. We have shown that with the protection of effective charge conjugation symmetry $\mathcal{C}$, zero modes can appear at high symmetry points. Since the charge conjugation transform a particle into its anti-particle, the quasiparticle corresponding to the $\mathcal{C}$-protected zero modes is its own anti-particle and can be identified as a Majorana quasiparticle.
@@
We have also shown that when the effective charge conjugation symmetry $\mathcal{C}$ is violated, bulk zero modes can also appear due to quantum numbers protected level crossing or $\mathcal{I}^{\prime}\mathcal{T}^{\prime}$ protected quantized Berry phase. Since the Hamiltonian is not commuting with $\mathcal{C}$, so the $\mathcal{C}$ transformed quasiparticle is not an eigenstate of the Hamiltonian. Therefore, strictly speaking, the zero modes without $\mathcal{C}$ symmetry is not a Majorana quasiparticle.
@@
Although $\mathcal{P}$ is always anti-commuting with the Hamiltonian, $\mathcal{P}$ transformed quasiparticles are not independent on the original quasiparticles since $\mathcal{P}$ is an redundancy of the SCs/SFs/QSLs.
@@
Classification of nodal SCs/SFs/QSLs. If two BdG systems have the same symmetry, and their low-energy quasiparticles are one-to-one corresponding to each other, then we can identify the two systems as a same phase. In this sense, we can provide a method to classify gapless nodal SCs/SFs/QSLs:
@@
(i) the symmetry group, which is an extension of the space group of the lattice by internal symmetries such as spin rotation and charge operations. A systematic study of all allowed symmetry groups for given lattice structure will be given elsewhereendnote72.
@@
(ii) the pattern of the bulk zero modes, whose positions form several sets of $k$ stars, namely $\{k^{\star}\}$, where each $k^{\star}$ is a set of symmetry related equivalent momentum points. For the set of $\{k^{\star}\}$, we also know the total number of zero energy nodes of the system.
@@
(iii) the degeneracy of the zero modes in every point of $\{k^{\star}\}$. The degeneracy of zero modes at equivalent $k$ points are the same.
@@
(iv) the dispersion and physical properties of the quasiparticles, which are described by the $k\cdot p$ effective theory.
@@
Effect of interactions. It can be expected that under weak interactions, the quasiparticles corresponding to the bulk zero modes remain gapless. However, with the increasing of the interaction strength, the fate of the quasiparticles is dependent on the degeneracy as well as the form of dispersion. Quantitative results should be obtained using renormalization group calculations, and will not be discussed here.
@@
In summary, using the theory of projective reps, we systematically studied the symmetry conditions of bulk zero modes in the mean field theory of superconductors, superfluids and $Z_{2}$ QSLs where fermions pair to form Cooper pairs. We than provided an efficient method to obtain the $k\cdot p$ effective theory of these gapless quasiparticles, from which one can get the dispersions as well as their physical response to external fields. The positions of the zero modes can be controlled by adjusting the symmetry of the Hamiltonian. Furthermore, the dispersion of the gapless quasiparticles can be linear or of higher order. The above results are illustrated via concrete lattice modes. Our symmetry representation-based theory is complementary to the topological origin (theory of symmetry indicators) of gapless quasiparticles in BdG systems, and is helpful for experimental realization of Majorana like gapless quasiparticles.
@@
We thank Yu-Xin Zhao, Zhi-Da Song, Chen Fang, Yuan-Ming Lu, Jiucai Wang and Bruce Normand for helpful discussions. A.Z.Y thank Xiao-Yue Wang for her helpful comments. This work is supported by NSFC (Grants No.12374166, 12134020) and National Basic Research and Development plan of China (Grants No.2023YFA1406500, 2022YFA1405300).
@@
$\tau_{x,y,z}$ generate the charge $SU(2)$ group and act on the particle-hole degrees of freedom. For instance the charge conjugation can be chosen as $\mathcal{C}=e^{-i{\tau_{x}\over 2}\pi}$. The product of $\tau_{x,y,z}$ with identity matrix are noted as $\Gamma_{x,y,z}=\tau_{x,y,z}\otimes I$.
$\sigma_{x,y,z}$ generate the spin $SU(2)$ group and act on the spin degrees of freedom.
@@
For an irrep $D(G)$ of anti-unitary group $G$ of the $\mathbb{C}$ or $\mathbb{H}$ type, one can reduce the restricted rep $D(H)$ into two irreps with $H$ the maximal unitary subgroup of $G$. $\omega_{x,y,z}$ acts on the irrep spaces of $D(H)$. The eigenvalue ±$\pm 1$ of $\omega_{z}$ respectively labels the two irrep spaces, and $\omega_{x,y}$ exchange the two subspaces.

### v3

@@ IV.2 Triangular lattice (spin-${1\over 2}$): $G_{b}=p3\times{Z}_{2}^{\mathcal{T}}$
one obtains the full symmetry group $F^{c}_{f}$. From the band representation, minimal $\mathcal{C}$-zero modes are found at the $\Gamma$ point and the three M points.
@@
$D(C_{3}^{z})=-I_{4},\ \ \ D(\mathcal{T})K=-\mathrm{i}I_{2}\otimes\sigma_{y}K,$
in the bases $\Psi^{\dagger}_{k}$. Without $\mathcal{C}$, the little co-group becomes $F_{f}(k)/Z_{2}^{f}=\mathscr{C}_{3}\times Z_{2}^{\mathcal{T}}\times Z_{2}^{\mathcal{P}}$. Defining $\tilde{\mathcal{P}}=\mathcal{T}\mathcal{P}$, then the above rep of $F_{f}(k)$ corresponds to a set of RMNZM of the case (6) in the Tab.1. When $\mathcal{C}$ is turned on, then the nonzero modes are turned into reducible minimal $\mathcal{C}$-zero modes. Furthermore, the $k\cdot p$ theory indicates that the lowest order dispersion is cubic (see Sec.III.1), and around the $\Gamma$ point, the third-order $k\cdot p$ Hamiltonian can be written as
@@
For the same reason as the $\Gamma$ point, the $\mathcal{C}$ symmetry turns the RMNZM into reducible minimal $\mathcal{C}$-zero modes. The $k\cdot p$ theory indicates that the energy dispersion is linear at the $M$ points (see Sec.III.1), for instance, the $k\cdot p$ Hamiltonian for the $(0,{1\over 2})$ point is given by
@@ IV.3 Triangular lattice (spinless): $G_{b}=p6m\times Z_{2}^{\mathcal{T}}$
The band representation of the above group $F^{c}_{f}$ indicates that the high symmetry lines on the mirror planes $\mathcal{M}_{1,2,3}$ (i.e. the lines linking $\Gamma$ and $K,K^{\prime}$ points) are nodal lines of irreducible $\mathcal{C}$-zero modes.
@@
(a) band structure with $\mathcal{M}^{\prime}_{1,2,3}$
(b) zero modes in (a)
(c) band structure without $\mathcal{M}^{\prime}_{1,2,3}$
(d) zero modes in (c)
@@ IV.4 Multi-node $Z_{2}$ QSL on Honeycomb lattice: $G_{b}=p\bar{3}1m\times Z_{2}^{\mathcal{T}}$
$\mathcal{T}^{\prime}=(P_{f}^{B}e^{-\mathrm{i}{\tilde{\tau}_{y}\over 2}{\pi}}e^{-\mathrm{i}{\tilde{\sigma}_{y}\over 2}{\pi}}K||\mathcal{T}),\ \mathcal{I}^{\prime}=(P_{f}^{B}||\mathcal{I}),$
@@
The particle-hole operation is simply represented as $\Gamma_{x}K$ as shown in (7).
@@
The spectrum of the Hamiltonian (31) can support a series of zero modes in the BZ and the corresponding state is a multi-node QSL. For instance, Fig.11 shows the band structure with parameters $\eta_{0}=2,\eta_{1}=1.2,\eta_{2}=0.8,\eta_{3}=0.5,\eta_{4}=0.2,\eta_{5}=1.2$ and $\eta^{\prime}_{0}=1.6,\eta^{\prime}_{1}=0.2,\eta^{\prime}_{2}=0.8,\eta^{\prime}_{3}=2.5,\eta^{\prime}_{4}=0,\eta^{\prime}_{5}=0$. There are totally $20$ cones, where each of the $2$ cones in green color contains irreducible zero modes protected by the little co-group $F_{f}(K)=\mathscr{C}^{\prime}_{3v}\times Z_{2}^{\mathcal{I}^{\prime}\mathcal{T}^{\prime}}$; the $6$ cones in blue color contain zero modes protected by crossing of energy bands carrying different quantum numbers of the mirror symmetries $\mathcal{M}^{\prime}_{1,2,3}$; the $12$ cones in red color contain zero modes protected by $\pi$-quantized Berry phase due to the $\mathcal{I^{\prime}T^{\prime}}$ symmetry.
@@
Figure 11: (a)The Band structure of the Hamiltonian (31) with parameters given in the maintext. The number of the cones is ‘$12+6+2$’, where the ‘$12$’ cones in red are due to zero modes protected by the $\mathcal{I^{\prime}T^{\prime}}$ symmetry, the ‘$6$’ cones in blue are protected by the mirror symmetries $\mathcal{M}^{\prime}_{1,2,3}$, and each of the ‘$2$’ green cones results from irreducible zero modes; (b) the dots illustrate the positions of the zero modes in (a).
The two high symmetry points $K=({1\over 3},{2\over 3})$ and $K^{\prime}=({2\over 3},{1\over 3})$ have the little co-group $F_{f}(k)/Z_{2}^{f}=\mathscr{C}^{\prime}_{3v}\times Z_{2}^{\mathcal{I^{\prime}T^{\prime}}}\times Z_{2}^{\mathcal{PT^{\prime}}}$. All of the $C^{\prime}_{3c}$, $\mathcal{M^{\prime}}_{1,2,3}$ symmetries can be treated as effective charge conjugation, and the 2-fold irreducible zero modes are represented as
@@
According to Eq.(III.1) one has $a_{({\rm v\times ph})}=1$. So the dispersion around the $K$ and $K^{\prime}$ points is linear and the $k\cdot p$ Hamiltonian has one free parameter $a$, yielding
$H_{\rm eff1}=a(\delta k_{x}\mu_{x}+\delta k_{y}\mu_{y}).$
Similarly, the $k\cdot p$ Hamiltonian for the cones at the high symmetry lines with mirror symmetries $\mathcal{M}^{\prime}_{1,2,3}$ have two free parameters $b,c$. For example, around the left blue cone, the effective Hamiltonian reads
$H_{\rm eff2}=$ $b\delta k_{x}{\sqrt{2}\over 2}(\cos\theta_{2}\mu_{x}-\sin\theta_{2}\mu_{y})$
$+c\delta k_{y}{\sqrt{2}\over 2}(\sin\theta_{2}\mu_{x}+\cos\theta_{2}\mu_{y})$
@@
$H_{\rm eff3}=(d\mu_{x}+e\mu_{y})\delta k_{x}+(f\mu_{x}+g\mu_{y})\delta k_{y},$ (32)
which has four parameters $d,e,f,g$.
@@ V Discussion and Conclusion
To be Majorana or not. We have shown that with the protection of effective charge conjugation symmetry $\mathcal{C}$, zero modes can appear at high-symmetry points. Since charge conjugation transforms a particle into its antiparticle, the quasiparticle corresponding to the $\mathcal{C}$-zero modes is its own antiparticle and can be identified as a Majorana quasiparticle.
@@
We have also shown that when effective charge conjugation symmetry $\mathcal{C}$ is violated, bulk zero modes can still appear due to quantum number-protected level crossings or $\mathcal{I}^{\prime}\mathcal{T}^{\prime}$-protected quantized Berry phases. Since the Hamiltonian does not commute with $\mathcal{C}$, the $\mathcal{C}$-transformed quasiparticle is not an eigenstate of the Hamiltonian. Therefore, strictly speaking, quasiparticles associated with zero modes without $\mathcal{C}$ symmetry are not of Majorana type.
@@
Although $\mathcal{P}$ always anti-commutes with the Hamiltonian, $\mathcal{P}$-transformed quasiparticles are not independent of the original quasiparticles since $\mathcal{P}$ is a redundancy of the SCs/SFs/QSLs.
@@
Classification of nodal SCs/SFs/QSLs. If two BdG systems have the same symmetry and their low-energy quasiparticles correspond one-to-one, we can identify the two systems as the same phase. In this sense, we propose a method to classify gapless nodal SCs/SFs/QSLs:
@@
(i) The symmetry group, which extends the space group of the lattice by internal symmetries such as spin rotation and charge operations.
@@
(ii) The pattern of the bulk zero modes, whose positions form several sets of $k$ stars, denoted $\{k^{\star}\}$, where each $k^{\star}$ is a set of symmetry-related equivalent momentum points. For each set $\{k^{\star}\}$, we determine the total number of zero-energy nodes in the system.
@@
(iii) The degeneracy of the zero modes at each point in $\{k^{\star}\}$. The degeneracy of zero modes at equivalent $k$ points is the same.
@@
(iv) The dispersion and physical properties of the quasiparticles, described by the $k\cdot p$ effective theory.
@@
Effect of interactions. It is expected that under weak interactions, the quasiparticles corresponding to the bulk zero modes remain gapless. However, with increasing interaction strength, the fate of the quasiparticles depends on the degeneracy and the form of dispersion. Quantitative results should be obtained using renormalization group calculations, which are beyond the scope of this work.
@@
In summary, using projective representation theory, we systematically studied the symmetry conditions for the existence of bulk zero modes in the mean-field theory of superconductors, superfluids, and $Z_{2}$ QSLs, where fermions pair to form Cooper pairs. We then provided an efficient method to obtain the $k\cdot p$ effective theory of these gapless quasiparticles, from which one can determine the dispersions and their physical responses to external fields. The positions of the zero modes can be controlled by adjusting the symmetry of the Hamiltonian. Furthermore, the dispersion of the gapless quasiparticles can be linear or of higher order. These results are illustrated with concrete lattice models. Our symmetry representation-based theory complements the topological origin (theory of symmetry indicators) of gapless quasiparticles in BdG systems and aids in the experimental realization of Majorana-like gapless quasiparticles.
@@
We thank Yu-Xin Zhao, Zhi-Da Song, Chen Fang, Yuan-Ming Lu, Jiucai Wang, Houmin Du, Zhen Feng and Bruce Normand for helpful discussions. A.Z.Y thank Xiao-Yue Wang for her helpful comments. This work is supported by NSFC (Grants No.12374166, 12134020) and National Key Research and Development Program of China (Grants No.2023YFA1406500, 2022YFA1405300).
@@
$\tau_{x,y,z}$ generate the charge $SU(2)_{c}$ group and act on the particle-hole degrees of freedom. For instance the charge conjugation can be chosen as $\mathcal{C}=e^{-i{\tau_{x}\over 2}\pi}$. The product of $\tau_{x,y,z}$ with identity matrix are noted as $\Gamma_{x,y,z}=\tau_{x,y,z}\otimes I$.
$\sigma_{x,y,z}$ generate the spin $SU(2)_{s}$ group and act on the spin degrees of freedom.
@@
For an irrep $D(G)$ of anti-unitary group $G$ of the $\mathbb{C}$ or $\mathbb{H}$ type, one can reduce the restricted rep $D(H)$ into two irreps with $H$ the maximal unitary subgroup of $G$. $\omega_{x,y,z}$ acts on the irrep spaces of $D(H)$. The eigenvalue ±$\pm 1$ of $\omega_{z}$ respectively labels the two irrep spaces, and $\omega_{x,y}$ exchange the two subspaces. Furthermore, the notation $\Omega_{x,y,z}=\omega_{x,y,z}\otimes I$ has been used in the main text.